Question:
Payments of $670 are being made at the end of each month for 5 years at an interest rate of 8% compounded monthly. Calculate the Present value?
1
Expert's answer
2019-12-03T10:04:26-0500

PV=PMT*(1 - 1/((1+i/k)n*k)/(i/k)

PV=PMT*(1 - 1/((1+0.08/12)5*12)/(0.08/12)

PMT = 670$

PV=670*(1-1/(1.0066)60)/0.0066

PV=670*(1-1/1.4839)/0.0066

PV=670*(1-0.6738)/0.0066

PV=670*0.3261/0.0066

PV=33104.09